Tom is older than Fred. The sum of their ages is 44. In 4 years Tom will be 20 years less than twice Freds age. Find their current ages.
Answer by htmentor(1343)

The sum of their current ages is 44: T + F = 44 -> T = 44 - F
In 4 years, Tom will be 20 years less than twice Fred's age:
T + 4 = 2(F + 4) - 20
44 - F + 4 = 2F - 12
3F = 60 -> F = 20
Fred is 20, and Tom is 24